A rocket of mass 1000kg moves vertically upward with an acceleration of 4.0m/s2 . What is the total force generated by it engines? (g=9.8m/s2 .)
1
Expert's answer
2021-08-09T16:24:36-0400

We can find the total force from the Newton's Second Law of Motion:


"F_{tot}-mg=ma,""F_{tot}=m(g+a),""F_{tot}=1000\ kg\cdot(9.8\ \dfrac{m}{s^2}+4.0\ \dfrac{m}{s^2})=13800\ N."